OpenStudy (anonymous):

I need some serious help with standard form, Im really confused . Here's a problem , Write the equation y - 2 = 2(x - 3) in standard form. Can someone help me to solve it so I can know how to solve other problems like it ?

OpenStudy (anonymous):

Standard form is \[y = mx + b\]Let's use our algebra to solve this. First, add 2 to both sides. \[y = 2(x-3) + 2\]Then distribute the 2, and combine like terms. \[y = 2x - 6 + 2 = 2x - 4\]

OpenStudy (anonymous):

Standard form is ax+by=c By solving above the equation we y-2=2x-6 2x-y=6-2 2x-y=4 is the standard form equation y=mx+c is not standard form but it is the slope-intercept form

OpenStudy (anonymous):

@Atchyut , do those steps apply to all standard form problems like this one?

OpenStudy (anonymous):

whoops! My form was slope-intercept. The same principles of algebra apply regardless of the required form.

OpenStudy (anonymous):

standard form is \[Ax+By=C\]

OpenStudy (anonymous):

y - 2 = 2(x - 3) The equation above is what is known as point slop form. There are three forms to an equation of a line and they are standard form, point slope form and slope intercept form. Standard form = Ax + By = C Point Slope = y - y1 = m(x - x1) Slope Intercept = y = mx + b In order to convert Point Slope or Slope Intercept from to standard form you have to get all the variables/terms to the left side and all the constants to the right. The C represents a constants. Ok lets look at your equation which is in Point Slope form. y - 2 = 2(x - 3) First we need to get rid of the parenthesis, which is 2(x - 3), and we do this by distributing the 2 y - 2 = 2x -6 Now we move the variables/terms to the left side. -2x + y - 2 = -6 Now we move our constants to the right side -2x + y = -6 + 2 Now we do the math on the right side, which is -6 + 2, -2x + y = -4 The above is in standard form now.

OpenStudy (anonymous):

Yeah this is the actual procedure and u can follow DoodleTech gave very good info about basics of straight lines above